OpenStudy (anonymous):

Question: "32 is 40% of what number?" >I'd attempted to work this out and got an outrageously high number! >I thought to multiply the two numbers...would I be dividing instead since it's to FIND the number the % is being taken from?? Thanks!

OpenStudy (johnweldon1993):

That would be correct...division \[\frac{ 32 }{ .40 } =?\]

OpenStudy (anonymous):

DANG IT! .....Thanks! should've went with my gut.... Ill see if I get the right answer now. Its SUPPOSED to be 80, so I'll see.

OpenStudy (anonymous):

\[\frac{ 40 }{100} *x = 32\\\frac{ 2 }{ 5} *x = 32\\x = 80\\ \]

OpenStudy (anonymous):

Thanks @Luis_Rivera !! Exactly what I needed! :)

OpenStudy (anonymous):

its quite simple just take 32 as 40% then 100% of the no. will be 32/40 *100 = 80

OpenStudy (anonymous):

" dying aint much of a living son "

OpenStudy (anonymous):

its the result of the simplification of 40/ 100

OpenStudy (anonymous):

divided by 20... but how'd you get from the second to last part to your answer? (sorry if I'm a hassle!) an thanks for your help!